The perimeter of a rectangle is 66 inches. The width of the rectangle is 11 inches. what is the rectangles ares
Answer by mananth(16946)

perimeter = 66
length =11
66= 2*(L+11)
33=L+11
L=22
...
Area =22*11= 242 sq.inches
m.ananth@hotmail.ca